The first week of the on-sky engineering campaign with LSSTCam followed a full schedule of daytime, evening, and nighttime activities, and saw rapid progress on many aspects of system commissioning.
One of the major accomplishments was a first demonstration of stable survey-mode operation of the active optics system (AOS). The AOS utilizes two control systems: (1) an open-loop system that determines corrections to the rigid body and mirror bending mode degrees of freedom based on a set of predefined look-up tables and (2) a closed-loop system that computes additional refinements based on near-realtime wavefront measurements acquired during operation. Both systems are needed to control the optical surfaces at submicron-scale precision and thereby maintain stable image quality as the telescope slews in elevation and azimuth, as the Camera rotates on its axis and changes filters, and as temperatures vary throughout the night.
As previously reported, during the ComCam on-sky engineering campaign, the team tested the closed loop system by pistoning the entire ComCam focal plane to measure the wavefront from intra- and extra-focal “donut” images of the telescope pupil. In practice, this involved slewing to a specific target, running a series of such closed-loop iterations to bring the system into focus and alignment, and then relying on the open loop while tracking the target for a sequence of observations for Science Pipelines commissioning. One of the many benefits of LSSTCam is that it comes equipped with pairs of intra- and extra-focal wavefront sensors on each of the 4 corner rafts so that the data for AOS closed-loop analysis can be taken at the same time as in-focus science observations. The team began by running the AOS closed loop using wavefront sensor measurements with repeated observations at a fixed telescope orientation to verify convergence of the optical state. By the end of the week, the team utilized the feature-based scheduler (FBS) to drive survey-mode observations including translational and rotational dithers and demonstrated the ability to refine the optical state continuously while maintaining stable image quality across the focal plane. Already on the third night of the campaign, the Observatory achieved a delivered image quality of 0.7” - 0.8” across the 3.5 deg diameter field of view, consistent with the required system contribution of ~0.4”. AOS commissioning will continue to be a major emphasis of the coming weeks as the team refines the pointing model and look-up tables for the open loop, measures the response to perturbing degrees of freedom, characterizes the reference wavefront including effects of intrinsic optical aberrations, and tunes the closed-loop system for a variety of observing conditions. These activities are intended to further reduce the ellipticity and to increase the uniformity of delivered image quality across the field of view.
The week also included extensive testing of the Camera filter exchange system. By the end of the week, the team demonstrated filter changes in the survey-mode operation described above, and acquired on-sky observations in the ugri bands within the same night. In each band, the AOS system typically converged within 3-4 iterations and was stable over many iterations. The delivered image quality was as expected in all bands, though limited on some nights by atmospheric seeing that was >1”. These observations enable early tests of the astrometric and photometric calibration, coaddition, and difference image analysis steps of the Science Pipelines.
Owing partly to the larger focal plane mosaic, the volume of on-sky pixel data collected during the first 3 nights of LSSTCam on-sky engineering exceeded that of the seven-week ComCam on-sky campaign.
In addition to the on-sky testing, the team is commissioning in-dome calibration systems with LSSTCam, including the calibration screen and associated light sources. The first task was an initial alignment of the broadband LED projector, reflector optic, telescope, and calibration screen. Initial sets of LED flats were acquired in the ugri bands.
LSSTCam and its support systems continued stable operation throughout the week. Three of the CCDs mentioned last week are again fully operational, with work continuing on the remaining three in parallel.
—
Want to help share these updates with your communities on social media? Make sure you’re following Rubin Observatory and the #CommissioningNotes hashtag — especially on X, BlueSky, Mastodon, and LinkedIn, where we post weekly simplified summaries of these updates!